# Towards Backdoor Stealthiness in Model Parameter Space

### v2

@@ 4.1. Experimental Setup
Datasets and Architectures. We follow the common settings in existing backdoor attacks and defenses and conduct experiments on CIFAR10 (Krizhevsky et al., 2009), GTSRB (Stallkamp et al., 2012), and a subset of ImageNet (Deng et al., 2009) with 200 classes and 1,300 images per class (ImageNet200). More details about the datasets can be found in App. A.1. The primary evaluation is performed using ResNet18 (He et al., 2016). Moreover, we evaluate Grond using four additional architectures, VGG16 (Simonyan and Zisserman, 2015), DenseNet121 (Huang et al., 2017), EfficientNet-B0 (Tan and Le, 2019), and one recent architecture InceptionNeXt (Yu et al., 2024) (see Table 15 in App. C.2). We also evaluate Grond with large and transformer-based models (see Table 16 in App. C.3).
Attack Baselines. Grond is compared with 12 representative attacks: BadNets (Gu et al., 2019), Blend (Chen et al., 2017), WaNet (Nguyen and Tran, 2021), IAD (Nguyen and Tran, 2020), AdvDoor (Zhang et al., 2021), BppAttack (Wang et al., 2022b), LC (Turner et al., 2019), Narcissus (Zeng et al., 2023), Adap-Blend (Qi et al., 2023a), SSDT (Mo et al., 2024), DFST (Cheng et al., 2021), and DFBA (Cao et al., 2024). The default poisoning rate is set at 5% (of the training set) for all attacks following previous work (Xu et al., 2024a, b). Additionally, Grond is evaluated under various poisoning rates to provide a thorough analysis of its effectiveness. Following related works, the training schedule for attacks is 200 epochs when using CIFAR10 and GTSRB, and 100 epochs for ImageNet200. We use 1,000 images as the validation set to select the best-performing checkpoint. More implementation details are provided in App. A.2.
Defense Baselines. We evaluate Grond and baseline attacks with 17 defenses, including four pruning-based methods (FP (Liu et al., 2018a), ANP (Wu and Wang, 2021), CLP (Zheng et al., 2022), and RNP (Li et al., 2023a)), five fine-tuning-based methods (vanilla FT, FT-SAM (Zhu et al., 2023), I-BAU (Zeng et al., 2022), FST (Min et al., 2023), and BTI-DBF (U) (Xu et al., 2024a)), five backdoor model detections (NC (Wang et al., 2019), Tabor (Guo et al., 2020), FeatureRE (Wang et al., 2022a), Unicorn (Wang et al., 2023), and BTI-DBF (Xu et al., 2024a)), two backdoor input detections (Scale-up (Guo et al., 2023) and IBD-PSC (Hou et al., 2024)), and a proactive defense CT (Qi et al., 2023b). Following their default settings, BTI-DBF (Xu et al., 2024a) and FP (Liu et al., 2018a) use 5% of training data, and other defenses use 1% of training data for detection or mitigation. CLP is a data-free backdoor pruning tool that uses no clean data. CT has access to the complete training set without knowing which samples are poisoned and can also interact with the model during training. Backdoor defense details and hyperparameters can be found in App. A.3.
@@ 4.2. Main Results on Backdoor Mitigation
Fine-tuning-based mitigations. Table 3 presents the backdoor performance against five fine-tuning-based defenses. In general, fine-tuning-based defenses are more effective than pruning-based defenses. For example, Narcissus and Adap-Blend can achieve ASRs higher than 60% against three out of four pruning-based defenses but are much less effective against most fine-tuning-based methods. FT-SAM is the most effective across all defenses, as shown in Tables 2 and 3, being able to compromise the effectiveness of all attack baselines. One important reason is that FT-SAM adopts Sharpness-Aware Minimization (Foret et al., 2021) to adjust the outlier of weight norm (large norms) to remove the potential backdoor. Larger weights of neurons are introduced by existing attacks to guarantee a high ASR (Liu et al., 2019), which also causes large differences when receiving benign and backdoor inputs (see Figure 4). Grond can bypass FT-SAM, as expected, since it deliberately decreases the weights of backdoor neurons, compromising the core working mechanism of FT-SAM.
@@
Comparison with Supply-Chain Attacks. Sharing a similar threat model to supply-chain attacks, we compare Grond and three state-of-the-art supply-chain attacks (SSDT (Mo et al., 2024), DFST (Cheng et al., 2021), and DFBA (Cao et al., 2024)), where these attacks are also designed to be robust against backdoor defenses. In particular, DFST (Cheng et al., 2021) proposes to include a controlled detoxification technique in the training process, which restrains the model from picking up simple features. DFBA (Cao et al., 2024) directly modifies a few parameters of a classifier to inject a backdoor. SSDT (Mo et al., 2024) introduces additional terms in the loss for the Source-Specific and Dynamic-Triggers (i.e., SSDT) attack, which obscures the difference between normal samples and malicious samples. Tables 2 and 3 also include the performance of supply-chain attacks against pruning- and fine-tuning-based defenses. It is clear that existing backdoor defenses can defeat supply-chain attacks. The ASR of DFST (Cheng et al., 2021), DFBA (Cao et al., 2024), and SSDT (Mo et al., 2024) are decreased to less than 10% while the BA drop is less than 3%.
@@ 4.3. Adaptive Defenses
Neuron noise targeting ABI. Second, we consider adding noise to neurons of Grond models as an adaptive defense, as the ABI component involves operations on neuron weights. Specifically, we add noise to the weights and biases of batch normalization layers. The range of noise is limited by $[-\epsilon_{noise},\epsilon_{noise}]$. Table 5 provides the performance of Grond and the benign model under different noise levels. With increased noise, the benign accuracy of both benign and Grond models is decreased, but the ASR of Grond remains high. This result also supports the stealthiness of Grond in the parameter space.
@@ 4.4. Backdoor Analysis
Method Benign Grond
@@
Grond’s weights are more difficult to be modified by backdoor defense. This section analyzes the changes while applying pruning- and fine-tuning-based defenses to baseline attacks and Grond. Our goal is to demonstrate that the baseline attacks can be significantly affected by defenses, but Grond can resist these defenses. In Figure 5, we record the changes in the weights of 512 neurons (in layer 4 of ResNet18) after applying two types of defenses, FT-SAM (Zhu et al., 2023) and CLP (Zheng et al., 2022), as they both focus on the backdoor-related neurons. In contrast to BadNets, the weight changes in Grond model after FT-SAM fine-tuning are smaller. In the sorted changes, it is clear that there are a few neurons for BadNets that correspond to significant changes, which is not the case for Grond. In the second row of Figure 5, the CLP can find a few neurons relevant to backdoor but cannot find these for Grond. In addition, pruning-based methods (as shown by CLP results in Figure 5) only improve a few neurons, while fine-tuning methods can update all neurons. We conjecture this is why fine-tuning-based defenses perform better than pruning-based defenses in Tables 2 and 3. More results with other baseline attacks are provided in Figure 12 in App. C.9.
@@
NC (Wang et al., 2019) 5 25% 2 10% 1 5%
Tabor (Guo et al., 2020) 5 25% 2 10% 0 0%
FeatureRE (Wang et al., 2022a) 0 0% 0 0% 0 0%
Unicorn (Wang et al., 2023) 0 0% 0 0% 0 0%
BTI-DBF (Xu et al., 2024a) 3 15% 5 25% 3 15%
@@ 4.6. Backdoor Detection
Following previous works (Xu et al., 2024a, b), we choose five representative backdoor model detections for evaluation. We use 20 models for each poisoning rate with different random seeds. Then, we report the number of models detected as backdoor models out of the 20. Table 6 shows that all detections fall short when detecting Grond. In particular, NC (Wang et al., 2019), Tabor (Wang et al., 2019), and BTI-DBF (Xu et al., 2024a) can detect a small part of backdoored models, while FeatureRE (Wang et al., 2022a) and Unicorn cannot detect any of them. For featureRE (Wang et al., 2022a), we conjecture it is over-dependent on the separability in the feature space, but Grond does not rely on prominent backdoor features according to Figure 9 in the Appendix App. C.4. For Unicorn (Wang et al., 2023), the false positive rate is high, and it tends to report every class as the backdoor target, even on models trained with benign data only. Except for model detection, Grond can also bypass input-space detections as demonstrated in App. C.1.
@@
Strategy No Defense CLP (Zheng et al., 2022) FT-SAM (Zhu et al., 2023)
@@
Strategies for [Uncaptioned image] No Defense CLP (Zheng et al., 2022) FT-SAM (Zhu et al., 2023)
@@
Arch Method No Defense CLP (Zheng et al., 2022) FT-SAM (Zhu et al., 2023)
@@ 4.7. Ablation Study
Trigger generation. To explore the influence of trigger patterns, we employ and evaluate three types of triggers: random noise, PGD perturbation, and UPGD perturbation, using ResNet18 on CIFAR10. The random noise is sampled from a uniform distribution, and the PGD employs a projected gradient descent to generate sample-wise perturbations (Madry et al., 2018). The generation of UPGD is described in Algorithm 1. All three triggers are limited to $8/255$ ($l_{\infty}$ norm) for imperceptibility and use the same training settings described in Table 11 in App. A.2.
Table 7 shows that random noise is ineffective as a backdoor trigger due to low ASR, even if no defense is applied. The sample-wise PGD perturbation is more effective than random noise and shows (limited) robustness against CLP and FT-SAM. UPGD generates the most effective backdoor trigger with an ASR higher than 80% after CLP and FT-SAM, and we speculate that the reason is that UPGD exploits features from the target class, similar to Narcissus (Zeng et al., 2023).
Concerning exploiting features from the target class, we also explore using the natural image as the trigger, which directly contains the semantic information. Table 8 shows the performance when using an automobile image from the CIFAR10 dataset as a semantic trigger. Specifically, inspired by naturally occurring backdoor (Khaddaj et al., 2023), we design the semantic trigger by resizing the automobile image to $8\times 8$ and sticking it on a part of (PR=5%) the training images. We consider four types of strategies, including only poisoning the same class as the trigger in clean-label, poisoning a different class (class airplane) in clean-label, only poisoning the same class as the trigger in dirty-label, and poisoning a different class (class airplane) in dirty-label. In Table 8, the semantic trigger can be effective as a backdoor trigger and resist against the CLP pruning. However, the semantic trigger is not robust against the FT-SAM fine-tuning. The reason is that the semantic trigger cannot effectively represent the feature of a class. Conversely, adversarial perturbation (PGD, UPGD) acquires more representative and discriminative class features, since they can capture the correlations among different regions of the decision boundary and easily fool most of the inputs (Moosavi-Dezfooli et al., 2017).
@@
Table 10. Evaluation with the proactive defense, CT (Qi et al., 2023b), under different poisoning rates (PR).
@@ 5.1. Proactive Defense
Real-world powerful defenders could take more initiative by intervening proactively in the attack process and exploiting poisoned data. We evaluate Grond against the SOTA proactive defense, CT (Qi et al., 2023b), that detects poisoned samples in the training data. Specifically, CT considers data from the original poisoned training data as regular batches and introduces randomly labeled benign data as confusing batches. Then, CT performs normal supervised training on both regular and confusing batches to produce an inference model, aiming to corrupt benign semantic features and correlations with correct labels in the inference model by confusing batches. The backdoor effect remains in the inference model because there is no trigger information in the confusing batches, and correctly predicted samples by the inference model are recorded as poisoned.
Table 10 presents the detection results on two baseline attacks (BadNets (Gu et al., 2019) and Adap-patch (Qi et al., 2023a)) and Grond. CT is effective against the two baseline attacks, where most poisoned samples in the training set are detected with a relatively low false positive rate. However, CT is not capable of detecting Grond when the poisoning rate is lower than 5% due to a high false positive rate and low recall. To understand why CT is not effective against Grond, we recall that the main idea of CT is to corrupt benign semantic features and their correct label, but not corrupt backdoor semantic features. However, Grond utilizes the benign semantic features of the target class to generate UPGD perturbation as the trigger. CT’s mechanism also corrupts the backdoor features of Grond. Therefore, CT cannot effectively detect Grond poisoned samples.
@@ 5.2. Visualization
Grad-CAM cannot spot the trigger area of Grond. Grad-CAM (Selvaraju et al., 2017) was originally designed to visualize the network’s preference when taking an input image. In backdoor defense research, Grad-CAM is leveraged to highlight the important areas in order to detect the potential backdoor trigger area (Cheng et al., 2020). Figure 8 shows the activated area of a clean model and Grond backdoored model using Grad-CAM. The activated area of Grond backdoored model is indistinguishable from the clean model, so the Grad-CAM-based defense (Cheng et al., 2020) is also ineffective against Grond.
t-SNE visualization of feature space. Figure 7 shows the latent feature (feature space of the last convolutional layers) from Grond backdoor models with and without adversarial backdoor injection in 2-D space and other baseline attacks by t-SNE (Van der Maaten and Hinton, 2008). The poisoning rate for all is 0.5%. WaNet cannot achieve satisfactory ASR at this very low poisoning rate, so we use the default setting according to their open-source implementation. Specifically, we perform dimensionality reduction for the latent features by t-SNE. The model architecture is ResNet18 and trained on CIFAR10. Each class of samples forms a tight cluster, and Grond poisoned samples are better mixed with the target class samples when the model is trained with adversarial backdoor injection.

### v3

@@ 4.1. Experimental Setup
Datasets and Architectures. We follow the common settings in existing backdoor attacks and defenses and conduct experiments on CIFAR10 (krizhevsky2009cifar10), GTSRB (Stallkamp2012gtsrb), and a subset of ImageNet (jia2009imagenet) with 200 classes and 1,300 images per class (ImageNet200). More details about the datasets can be found in App. A.1. The primary evaluation is performed using ResNet18 (He_2016_resnet). Moreover, we evaluate Grond using four additional architectures, VGG16 (Karen_2015_vgg), DenseNet121 (huang2017densely), EfficientNet-B0 (tan2019efficientnet), and one recent architecture InceptionNeXt (Yu2024inceptionnext) (see Table 15 in App. C.2). We also evaluate Grond with large and transformer-based models (see Table 16 in App. C.3).
Attack Baselines. Grond is compared with 12 representative attacks: BadNets (Gu_2019_badnet), Blend (chen_2017_blend), WaNet (nguyen2021wanet), IAD (nguyen2020inputaware), AdvDoor (zhang2021advdoor), BppAttack (Wang_2022_bpp), LC (turner2019lc), Narcissus (zeng2023narcisus), Adap-Blend (qi2022revisitingadaptive), SSDT (mo2024ssdt), DFST (Cheng2021dfst), and DFBA (cao2024dfba). The default poisoning rate is set at 5% (of the training set) for all attacks following previous work (xu2024btidbf; xu2024ban). Additionally, Grond is evaluated under various poisoning rates to provide a thorough analysis of its effectiveness. Following related works, the training schedule for attacks is 200 epochs when using CIFAR10 and GTSRB, and 100 epochs for ImageNet200. We use 1,000 images as the validation set to select the best-performing checkpoint. More implementation details are provided in App. A.2.
Defense Baselines. We evaluate Grond and baseline attacks with 17 defenses, including four pruning-based methods (FP (liu2018finepruning), ANP (Wu_2021_anp), CLP (zheng2022clp), and RNP (li2023rnp)), five fine-tuning-based methods (vanilla FT, FT-SAM (Zhu2023ftsam), I-BAU (zeng2022ibau), FST (min2023fst), and BTI-DBF (U) (xu2024btidbf)), five backdoor model detections (NC (wang_2019_nc), Tabor (guo_2020_tabor), FeatureRE (wang2022featurere), Unicorn (wang2023unicorn), and BTI-DBF (xu2024btidbf)), two backdoor input detections (Scale-up (guo2023scaleup) and IBD-PSC (hou2024ibdpsc)), and a proactive defense CT (qi2023proactivedetection). Following their default settings, BTI-DBF (xu2024btidbf) and FP (liu2018finepruning) use 5% of training data, and other defenses use 1% of training data for detection or mitigation. CLP is a data-free backdoor pruning tool that uses no clean data. CT has access to the complete training set without knowing which samples are poisoned and can also interact with the model during training. Backdoor defense details and hyperparameters can be found in App. A.3.
@@ 4.2. Main Results on Backdoor Mitigation
Fine-tuning-based mitigations. Table 3 presents the backdoor performance against five fine-tuning-based defenses. In general, fine-tuning-based defenses are more effective than pruning-based defenses. For example, Narcissus and Adap-Blend can achieve ASRs higher than 60% against three out of four pruning-based defenses but are much less effective against most fine-tuning-based methods. FT-SAM is the most effective across all defenses, as shown in Tables 2 and 3, being able to compromise the effectiveness of all attack baselines. One important reason is that FT-SAM adopts Sharpness-Aware Minimization (foret2021sam) to adjust the outlier of weight norm (large norms) to remove the potential backdoor. Larger weights of neurons are introduced by existing attacks to guarantee a high ASR (liu_2019_abs), which also causes large differences when receiving benign and backdoor inputs (see Figure 4). Grond can bypass FT-SAM, as expected, since it deliberately decreases the weights of backdoor neurons, compromising the core working mechanism of FT-SAM.
@@
Comparison with Supply-Chain Attacks. Sharing a similar threat model to supply-chain attacks, we compare Grond and three state-of-the-art supply-chain attacks (SSDT (mo2024ssdt), DFST (Cheng2021dfst), and DFBA (cao2024dfba)), where these attacks are also designed to be robust against backdoor defenses. In particular, DFST (Cheng2021dfst) proposes to include a controlled detoxification technique in the training process, which restrains the model from picking up simple features. DFBA (cao2024dfba) directly modifies a few parameters of a classifier to inject a backdoor. SSDT (mo2024ssdt) introduces additional terms in the loss for the Source-Specific and Dynamic-Triggers (i.e., SSDT) attack, which obscures the difference between normal samples and malicious samples. Tables 2 and 3 also include the performance of supply-chain attacks against pruning- and fine-tuning-based defenses. It is clear that existing backdoor defenses can defeat supply-chain attacks. The ASR of DFST (Cheng2021dfst), DFBA (cao2024dfba), and SSDT (mo2024ssdt) are decreased to less than 10% while the BA drop is less than 3%.
@@ 4.3. Adaptive Defenses
Neuron noise targeting ABI. Second, we consider adding noise to neurons of Grond models as an adaptive defense, as the ABI component involves operations on neuron weights. Specifically, we add noise to the weights and biases of batch normalization layers. The range of noise is limited by $\epsilon_{noise},\epsilon_{noise}]$. Table 5 provides the performance of Grond and the benign model under different noise levels. With increased noise, the benign accuracy of both benign and Grond models is decreased, but the ASR of Grond remains high. This result also supports the stealthiness of Grond in the parameter space.
@@ 4.4. Backdoor Analysis
Method
Benign Grond
@@
Grond’s weights are more difficult to be modified by backdoor defense. This section analyzes the changes while applying pruning- and fine-tuning-based defenses to baseline attacks and Grond. Our goal is to demonstrate that the baseline attacks can be significantly affected by defenses, but Grond can resist these defenses. In Figure 5, we record the changes in the weights of 512 neurons (in layer 4 of ResNet18) after applying two types of defenses, FT-SAM (Zhu2023ftsam) and CLP (zheng2022clp), as they both focus on the backdoor-related neurons. In contrast to BadNets, the weight changes in Grond model after FT-SAM fine-tuning are smaller. In the sorted changes, it is clear that there are a few neurons for BadNets that correspond to significant changes, which is not the case for Grond. In the second row of Figure 5, the CLP can find a few neurons relevant to backdoor but cannot find these for Grond. In addition, pruning-based methods (as shown by CLP results in Figure 5) only improve a few neurons, while fine-tuning methods can update all neurons. We conjecture this is why fine-tuning-based defenses perform better than pruning-based defenses in Tables 2 and 3. More results with other baseline attacks are provided in Figure 12 in App. C.9.
@@
NC (wang_2019_nc) 5 25% 2 10% 1 5%
Tabor (guo_2020_tabor) 5 25% 2 10% 0 0%
FeatureRE (wang2022featurere) 0 0% 0 0% 0 0%
Unicorn (wang2023unicorn) 0 0% 0 0% 0 0%
BTI-DBF (xu2024btidbf) 3 15% 5 25% 3 15%
@@ 4.6. Backdoor Detection
Following previous works (xu2024btidbf; xu2024ban), we choose five representative backdoor model detections for evaluation. We use 20 models for each poisoning rate with different random seeds. Then, we report the number of models detected as backdoor models out of the 20. Table 6 shows that all detections fall short when detecting Grond. In particular, NC (wang_2019_nc), Tabor (wang_2019_nc), and BTI-DBF (xu2024btidbf) can detect a small part of backdoored models, while FeatureRE (wang2022featurere) and Unicorn cannot detect any of them. For featureRE (wang2022featurere), we conjecture it is over-dependent on the separability in the feature space, but Grond does not rely on prominent backdoor features according to Figure 9 in the Appendix App. C.4. For Unicorn (wang2023unicorn), the false positive rate is high, and it tends to report every class as the backdoor target, even on models trained with benign data only. Except for model detection, Grond can also bypass input-space detections as demonstrated in App. C.1.
@@
Strategy No Defense CLP (zheng2022clp) FT-SAM (Zhu2023ftsam)
@@
Strategies for [Uncaptioned image] No Defense CLP (zheng2022clp) FT-SAM (Zhu2023ftsam)
@@
Arch Method No Defense CLP (zheng2022clp) FT-SAM (Zhu2023ftsam)
@@ 4.7. Ablation Study
Trigger generation. To explore the influence of trigger patterns, we employ and evaluate three types of triggers: random noise, PGD perturbation, and UPGD perturbation, using ResNet18 on CIFAR10. The random noise is sampled from a uniform distribution, and the PGD employs a projected gradient descent to generate sample-wise perturbations (madry2018pgd). The generation of UPGD is described in Algorithm 1. All three triggers are limited to $8/255$ ($l_{\infty}$ norm) for imperceptibility and use the same training settings described in Table 11 in App. A.2.
Table 7 shows that random noise is ineffective as a backdoor trigger due to low ASR, even if no defense is applied. The sample-wise PGD perturbation is more effective than random noise and shows (limited) robustness against CLP and FT-SAM. UPGD generates the most effective backdoor trigger with an ASR higher than 80% after CLP and FT-SAM, and we speculate that the reason is that UPGD exploits features from the target class, similar to Narcissus (zeng2023narcisus).
Concerning exploiting features from the target class, we also explore using the natural image as the trigger, which directly contains the semantic information. Table 8 shows the performance when using an automobile image from the CIFAR10 dataset as a semantic trigger. Specifically, inspired by naturally occurring backdoor (khaddaj2023rethinking), we design the semantic trigger by resizing the automobile image to $8\times 8$ and sticking it on a part of (PR=5%) the training images. We consider four types of strategies, including only poisoning the same class as the trigger in clean-label, poisoning a different class (class airplane) in clean-label, only poisoning the same class as the trigger in dirty-label, and poisoning a different class (class airplane) in dirty-label. In Table 8, the semantic trigger can be effective as a backdoor trigger and resist against the CLP pruning. However, the semantic trigger is not robust against the FT-SAM fine-tuning. The reason is that the semantic trigger cannot effectively represent the feature of a class. Conversely, adversarial perturbation (PGD, UPGD) acquires more representative and discriminative class features, since they can capture the correlations among different regions of the decision boundary and easily fool most of the inputs (Moosavi2017uap).
@@
Table 10. Evaluation with the proactive defense, CT (qi2023proactivedetection), under different poisoning rates (PR).
@@ 5.1. Proactive Defense
Real-world powerful defenders could take more initiative by intervening proactively in the attack process and exploiting poisoned data. We evaluate Grond against the SOTA proactive defense, CT (qi2023proactivedetection), that detects poisoned samples in the training data. Specifically, CT considers data from the original poisoned training data as regular batches and introduces randomly labeled benign data as confusing batches. Then, CT performs normal supervised training on both regular and confusing batches to produce an inference model, aiming to corrupt benign semantic features and correlations with correct labels in the inference model by confusing batches. The backdoor effect remains in the inference model because there is no trigger information in the confusing batches, and correctly predicted samples by the inference model are recorded as poisoned.
Table 10 presents the detection results on two baseline attacks (BadNets (Gu_2019_badnet) and Adap-patch (qi2022revisitingadaptive)) and Grond. CT is effective against the two baseline attacks, where most poisoned samples in the training set are detected with a relatively low false positive rate. However, CT is not capable of detecting Grond when the poisoning rate is lower than 5% due to a high false positive rate and low recall. To understand why CT is not effective against Grond, we recall that the main idea of CT is to corrupt benign semantic features and their correct label, but not corrupt backdoor semantic features. However, Grond utilizes the benign semantic features of the target class to generate UPGD perturbation as the trigger. CT’s mechanism also corrupts the backdoor features of Grond. Therefore, CT cannot effectively detect Grond poisoned samples.
@@ 5.2. Visualization
Grad-CAM cannot spot the trigger area of Grond. Grad-CAM (Selvaraju2017gradcam) was originally designed to visualize the network’s preference when taking an input image. In backdoor defense research, Grad-CAM is leveraged to highlight the important areas in order to detect the potential backdoor trigger area (xu2020defendingbygradcam). Figure 8 shows the activated area of a clean model and Grond backdoored model using Grad-CAM. The activated area of Grond backdoored model is indistinguishable from the clean model, so the Grad-CAM-based defense (xu2020defendingbygradcam) is also ineffective against Grond.
t-SNE visualization of feature space. Figure 7 shows the latent feature (feature space of the last convolutional layers) from Grond backdoor models with and without adversarial backdoor injection in 2-D space and other baseline attacks by t-SNE (van2008tsne). The poisoning rate for all is 0.5%. WaNet cannot achieve satisfactory ASR at this very low poisoning rate, so we use the default setting according to their open-source implementation. Specifically, we perform dimensionality reduction for the latent features by t-SNE. The model architecture is ResNet18 and trained on CIFAR10. Each class of samples forms a tight cluster, and Grond poisoned samples are better mixed with the target class samples when the model is trained with adversarial backdoor injection.